• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 07:39:45 on September15,2009

    Entry number 289628

    keyword=Owl shift summary.

    1:30 Q1 is dwon. Talked to run coordinator. Followed the
    instructions on web, could not fix it.
    2:00 Page to tech on call twice.
    2:15 Beam has multiple drops.
    2:35 Beam current is slowly ramping up to 100 uA. Now its 70uA.
    2:50 Current is 94uA.
    3:05 Heater power is low.
    4:00 Page jack to fix Q1, jack replied.
    4:40 MCC cannot give us 100uA, current is 94uA.
    4:43 Jack is here and a short controlled access is requested to fix Q1

    4:50 We do not have 100 uA for last two hours, we had 94uA.
    5:40 Ch 1 and ch 7 are zero.
    7:00 No beam.